OpenStudy (anonymous):

help in this question. AG3 hope you guys understand (x-4)/(x^2-2x-8)-(x-2)/(x2-16)

OpenStudy (anonymous):

\[\frac{x-4}{(x-4)(x+2)}-\frac{x-2}{(x-4)(x+4)}\] LCD of (x-4)(x+4)(x+2) \[(x-4)(x+4)(x-2)\frac{x-4}{(x-4)(x+2)}-(x-4)(x+4)(x+2)\frac{x-2}{(x-4)(x+4)}\] \[\frac{(x-4)(x+4)-(x-2)(x+2)}{(x-4)(x+4)(x+2)}\] \[\frac{x^2-16-x^2+4}{(x-4)(x+4)(x+2)}=\frac{-12}{(x-4)(x+4)(x+2)}\]

OpenStudy (anonymous):

actually shouldve cancelled the x-4's, oh well

OpenStudy (anonymous):

thanks man, Does not matter if is wrong, really thanks.

OpenStudy (anonymous):

its same answer, would've saved me some steps though